# Generic graded contractions of Lie algebras

### v1

@@ 1 Introduction
Deformation theory of different algebraic and geometric objects has been a subject of significant interest in mathematics ever since the study of Riemann surfaces and their moduli spaces by Riemann in XIX century, see a review of the history of the subject in [15]. There are various types of deformations appearing in algebra: one-parameter deformations, degenerations and contractions. From the perspective of physics, the most useful deformations are those that connect symmetry groups (or their Lie algebras) when one physical theory can be obtained as a limit of another. For example, Poincaré group, which is the symmetry group of relativistic mechanics, ‘transforms’ into Galilean group, the symmetry group of classical mechanics, when the speed of light is assumed to approach infinity. This example motivated Segal [35] to initiate the study of contractions of Lie groups and Lie algebras. Later, Inönü and Wigner [22, 23] considered the simplest types of one-parameter continuous contractions, where some of the basis elements of the Lie algebra are scaled by a parameter that tends to zero. This type of contractions is now called the Inönü–Wigner contractions. For example, the transition from Heisenberg algebra to abelian algebra of the same dimension, which represents the transition from quantum mechanics to classical mechanics when Planck’s constant goes to zero, is an Inönü–Wigner contraction. Another famous case of an Inönü–Wigner contraction is the transition from de Sitter invariant relativity to special relativity, corresponding to a contraction from de Sitter algebra to Poincaré algebra.
@@
Despite their simplicity, Inönü–Wigner contractions found a wide range of applications in both mathematics and physics. Saletan [34] and then Doebner and Melsheimer [10] generalized this concept in several directions, leading to what are now known as generalized Inönü–Wigner contractions. These generalized contractions provide a more convenient framework for applications to the study of representations, invariants and special functions, see, for example, [37]. For a more comprehensive literature review on this topic, as well as the state of the art on contractions between low-dimensional real and complex Lie algebras, see, e.g., [31]. Contractions of some infinite-dimensional Lie algebras (Witt, Virasoro and affine Kac–Moody algebras) were studied in [16, 17].
Contractions of Lie algebras are closely related to degenerations. In fact, over the field of complex numbers these two concepts coincide due to Borel’s closed orbit lemma. Milnor, in his seminal work on invariant metrics on Lie groups [26], implicitely used degenerations of Lie algebras to demonstrate the existence of invariant Rimannian metrics on Lie groups.
@@
Essentially, degenerations of an $n$-dimensional Lie algebra $\mathfrak{L}$ are limit points of its ${\rm GL}_{n}(\mathbb{F})$-orbit $\mathcal{O}(\mathfrak{L})$ in the variety of $n$-dimensional Lie algebras endowed with the Zariski topology. When the orbit $\mathcal{O}(\mathfrak{L})$ is a Zariski open set, the Lie algebra $\mathfrak{L}$ is called rigid. Intuitively, a Lie algebra is rigid if all Lie algebras close to it (in the sense of Zariski topology) are isomorphic. Rigidity of Lie algebras has been extensively studied and characterized in different terms, see, e.g., [5, 18] and references therein. The classification of rigid Lie algebras remains an important open problem.
@@
Over the past three decades, graded contractions have been extensively studied from various perspectives, including computations for specific graded Lie algebras [3, 6, 8, 9, 20, 21, 25, 29] and theoretical aspects [28, 39, 40]. The concept of graded contraction is flexible with respect to changes in the underlying category of algebras and thus can be easily transfered to other structures, such as superalgebras [30] and Jordan algebras [24]. For a review of the state of the art in graded contractions, we refer to the recent paper by Draper, Sánchez-Ortega and Meyer [9].
@@
For generic graded contractions, which are studied in [28, 37, 39] and are also the subject of the present paper, the defining conditions are chosen in such a way that the contraction can be applied uniformly to all $G$-graded Lie algebras. This allows us to study such contractions from various perspectives: group cohomology, affine algebraic geometry and monoidal categories. As a result, we not only generalize and reinterpret existing results on classification of such contractions from [28, 39], but also place them into a wider theoretical framework.
@@
The structure of the paper is as follows. In Section 2, theoretical background on gradings, contractions, degenerations, and graded contractions as well as their equivalences are reviewed.
@@
In Section 3, we discuss the classification of generic graded contractions. To any subset $S$ of the set of unordered pairs in $G$, we associate free abelian groups of “surviving defining relations” and “surviving higher-order identities” and use them to show that contractions with support $S$ are classified by a certain abelian group $H^{2}_{S}(\mathbb{F}^{\times})$, which can be computed using the exact sequences presented in Theorem 19. In particular, for an algebraically closed or a real closed field $\mathbb{F}$, we reinterpret the results of Weimar–Woods [39].
@@
Section 4 is devoted to the study of the affine variety of generic graded contractions. We study the action of an algebraic torus on this variety and the orbit closures of contractions. In Theorem 27, we describe the vanishing ideal of such an orbit closure. We show that the contractions in the closure of the unique open orbit are precisely the “continuous contractions” in the sense of [28] and relate them to Inönü–Wigner contractions using Theorem 19 and [38, Theorem VI.2].
@@
In Section 5, we discuss the Weimar–Woods conjecture that two graded contractions are strongly equivalent if and only if they are equivalent via normalization. We show that a generic $G$-graded contraction endows the identity endofunctor of the monoidal category of $G$-graded vector spaces with a lax monoidal structure. This allows us to prove a functorial version of the Weimar–Woods conjecture: two graded contractions define naturally isomorphic lax monoidal endofunctors if and only if they are equivalent via normalization (Proposition 36). In Section 6, we summarize the results of the paper, discuss open problems and possible directions for future research.
@@ 2 Preliminaries
We review some basic facts about contractions, gradings and graded contractions of Lie algebras following [9, 13, 30, 31].
@@ 2.1 Contractions and degenerations of Lie algebras
Contractions. Let $\mathfrak{L}:=(V,[\cdot,\cdot])$ be a real or complex Lie algebra with an underlying vector space $V$ and a Lie bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ]. Consider a continuous function $U\colon(0,1]\to{\rm GL}(V)$. A parameterized family of new Lie brackets on $V$ is defined in terms of the Lie bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] in the following way:
$\forall t\in(0,1],\quad\forall x,y\in V\colon\qquad[x,y]_{t}:=U_{t}^{-1}[U_{t}x,U_{t}y].$ (1)
@@
If the limit $[x,y]_{0}:=\lim_{t\to 0^{+}}[x,y]_{t}$ exists for any $x,y\in V$, then $[\cdot,\cdot]_{0}$ is a well-defined Lie bracket on $V$ and the Lie algebra $\mathfrak{L}_{0}:=(V,[\cdot,\cdot]_{0})$ is called a one-parametric contionuous contraction (or simply contraction) of the Lie algebra $\mathfrak{L}=(V,[\cdot,\cdot])$.
@@
Choosing a basis in each subspace $V^{(s)}$, we represent $U_{t}$ by a diagonal matrix with entries $t^{n_{s}}$, each repeated $\dim V^{(s)}$ times. Then the limit $[\cdot,\cdot]_{0}$ exists if and only if the following constraints are satisfied:
@@
The constraints (3) are necessary and sufficient to ensure that $[\mathfrak{L}(n),\mathfrak{L}(m)]\subset\mathfrak{L}(n+m)$. The associated graded algebra $\operatorname{gr}\mathfrak{L}$ of $\mathfrak{L}$ with respect to the above filtration is isomorphic to the contracted algebra $\mathfrak{L}_{0}$ under the contraction matrix $U(\varepsilon)$ from (2). This follows directly from the description of the structure tensor $\tilde{c}_{ij}^{k}$ of $\mathfrak{L}_{0}$, see, e.g., [31, Section 3]:
@@
When $\mathbb{F}$ is algebraically closed, the orbit $\mathcal{O}(\mu)$ is open in its Zariski closure $\overline{\mathcal{O}(\mu)}$ by the closed orbit lemma in [4]. Hence, the boundary $\overline{\mathcal{O}(\mu)}\setminus\mathcal{O}(\mu)$ is a union of orbits of strictly lower dimension. Moreover, all points of $\mathcal{O}(\mu)$ are regular in $\overline{\mathcal{O}(\mu)}$. Finally, the orbit $\mathcal{O}(\mu)$ is a constructible set, so in the case $\mathbb{F}=\mathbb{C}$, the closure of $\mathcal{O}(\mu)$ in Zariski topology coincides with its closure in the metric topology, see [27, Corollary 1, p. 60]. Therefore, over $\mathbb{C}$, the concepts of contraction and degeneration coincide.
@@ 2. $(ii^{\prime})$
To prove the necessity, it suffices to find a $G$-graded Lie algebra $\mathfrak{L}$ satisfying (a) $[\mathfrak{L}_{g},\mathfrak{L}_{h}]\neq 0$ for all distinct $g,h\in G$, (b) for all $g,h,k\in G$ there exist $x\in\mathfrak{L}_{g}$, $y\in\mathfrak{L}_{h}$, $z\in\mathfrak{L}_{k}$ such that $[x,[y,z]]$ and $[y,[z,x]]$ are linearly independent. For instance, consider the free Lie algebra $\mathfrak{F}$ on the generators $x_{g}$, $y_{g}$ and $z_{g}$, where $g$ runs through $G$. This algebra has a natural grading by setting the $G$-degree of generators $x_{g}$, $y_{g}$ and $z_{g}$ to be $g$. Since $[x_{g},x_{h}]\neq 0$ for all $g\neq h$, condition $(i)$ implies $(i^{\prime})$ for the graded Lie algebra $\mathfrak{F}$. Since the commutators $[x_{g}[y_{h},z_{k}]]$ and $[y_{h}[z_{k},x_{g}]]$ are linearly independent in $\mathfrak{F}$, condition $(ii)$ implies $(ii^{\prime})$. ∎
@@
When $\varepsilon$ does not take value 00, conditions $(i^{\prime})$ and $(ii^{\prime})$ say that $\varepsilon$ is a symmetric two-cocycle with values in $\mathbb{F}^{\times}$ with trivial $G$-action: $\varepsilon\in Z^{2}_{sym}(G,\mathbb{F}^{\times})$.
@@
The obvious examples of the graded contractions are the constant maps $\varepsilon=1$ and $\varepsilon=0$. Their corresponding contracted algebras are respectively the original algebra $\mathfrak{L}$ and the abelian Lie algebra. Both these maps are examples of trivial graded contractions, that is, the graded contractions $\varepsilon\colon G\times G\to\mathbb{F}$ such that the for all $G$-graded Lie algebras $\mathfrak{L}$ the contracted algebra $\mathfrak{L}^{\varepsilon}$ is isomorphic to $\mathfrak{L}$ or to the abelian Lie algebra on the vector space $\mathfrak{L}$, see [39, Definition 2.6]
A more interesting example of a trivial graded contraction is that arising from the normalization of homogeneous components of the grading $\Gamma$. Let $\alpha\colon G\to\mathbb{F}^{\times}$ be an arbitrary map, then the map $d\alpha\colon G\times G\to\mathbb{F}^{\times}$,
$d\alpha(g,h)=\frac{\alpha(g)\alpha(h)}{\alpha(gh)},\quad\forall g,h\in G,$
is a graded contraction of $\mathfrak{L}$. It is clear that $\mathfrak{L}^{d\alpha}$ is isomorphic to $\mathfrak{L}$. The choice of the notation $d\alpha$ for this kind of contractions is justified by the fact that it is the coboundary of the map $\alpha$ if we consider $\mathbb{F}^{\times}$ as a $G$-module with trivial action. Furthermore, when $\alpha$ is a group homomorphism then $d\alpha=1$.
Graded degenerations. Let $G$ be an abelian group. Consider a $G$-grading $\Gamma$ on a finite-dimensional vector space $V$ over an algebraically closed field $\mathbb{F}$, $V=\bigoplus_{g\in G}V_{g}$, with $\dim V_{g}=n_{g}\in\mathbb{N}_{0}$. Denote $\underline{n}=(n_{g})_{g\in G}$ and let $\mathcal{L}_{\underline{n}}(\mathbb{F})$ be the variety of Lie algebra laws on $V$ that make it a $G$-graded algebra. The set $\mathcal{L}_{\underline{n}}(\mathbb{F})$ is an algebraic set in the affine space $\wedge^{2}V^{*}\otimes V$, which is invariant under the natural action of the subgroup $\prod_{g\in G}{\rm GL}(V_{g})$ of ${\rm GL}(V)$. We will consider the subgroup $D(\Gamma)$ of $\prod_{g\in G}{\rm GL}(V_{g})$ consisting of the operators that act as (nonzero) scalars on each $V_{g}$. It is clear that $D(\Gamma)$ is an algebraic torus of dimension $|{\rm Supp}(\Gamma)|$.
@@
Consider the free Lie algebra $\mathfrak{F}$ on the generators $x_{g}$ and $y_{g}$, where $g$ runs through $G$. We can define a $G$-grading on $\mathfrak{F}$ by setting the $G$-degree of generators $x_{g}$ and $y_{g}$ to be $g$. Given a set $S\subset\mathcal{P}_{G}$, consider the ideal $I_{S}$ of $\mathfrak{F}$ generated by $\mathfrak{F}^{3}$ (the third term of the lower central series of $\mathfrak{F}$) and the commutators $[\mathfrak{F}_{g},\mathfrak{F}_{h}]$ for all $\{g,h\}\notin S$. Then the quotient $\mathfrak{L}:=\mathfrak{F}/I_{S}$ is a $G$-graded Lie algebra, whose second support is $S$. Moreover, if $G$ is finite, the algebra $\mathfrak{L}$ is finite-dimensional. ∎
@@
Equivalences of graded contractions. For a specific fixed grading $\Gamma$ of a Lie algebra $\mathfrak{L}$, graded contractions are studied modulo several natural equivalences, see [9, Definitions 2.4, 2.6, 2.11]. We translate these notions into the generic case.

### v2

@@ 1 Introduction
Deformation theory of different algebraic and geometric objects has been a subject of significant interest in mathematics ever since the study of Riemann surfaces and their moduli spaces by Riemann in the XIX century, see a historical review of deformation theory in [19]. The inverse process to deformation is that of degeneration [32] and, in particular, contraction. A well-known example from physics is the contraction of Poincaré group of special relativity to Galilean group of classical mechanics when the speed of light is assumed to approach infinity. This example motivated Segal [43] to initiate the study of contractions of Lie groups and Lie algebras. Later, Inönü and Wigner [27, 28] considered the simplest types of one-parameter continuous contractions, where some of the basis elements of the Lie algebra are scaled by a parameter that tends to zero. This type of contractions is now called the Inönü–Wigner contractions. Saletan [42] and then Doebner and Melsheimer [13] generalized this concept in several directions, leading to what are now known as generalized Inönü–Wigner contractions.
@@
Contractions have found applications in the study of representations of Lie groups [14] and Lie algebras [6, 46], invariants [50] and special functions [29]. For a more comprehensive literature review on this topic, as well as the state of the art on contractions between low-dimensional real and complex Lie algebras, see, e.g., [39]. Contractions of some infinite-dimensional Lie algebras (Witt, Virasoro and affine Kac–Moody algebras) were studied in [20, 21].
@@
Contractions of Lie algebras are closely related to degenerations. In fact, over the field of complex numbers these two concepts coincide due to Borel’s closed orbit lemma. Milnor, in his seminal work on invariant metrics on Lie groups [34], used degenerations of Lie algebras to demonstrate the existence of invariant Riemannian metrics on Lie groups. Essentially, degenerations of an $n$-dimensional Lie algebra $\mathfrak{L}$ are limit points of its ${\rm GL}_{n}(\mathbb{F})$-orbit $\mathcal{O}(\mathfrak{L})$ in the variety of $n$-dimensional Lie algebras endowed with the Zariski topology. When the orbit $\mathcal{O}(\mathfrak{L})$ is a Zariski open set, the Lie algebra $\mathfrak{L}$ is called rigid. Intuitively, a Lie algebra is rigid if all Lie algebras close to it (in the sense of Zariski topology) are isomorphic. Rigidity of Lie algebras has been extensively studied and characterized in different terms, see, e.g., [5, 22] and references therein. The classification of rigid Lie algebras remains an important open problem.
@@
Over the past three decades, graded contractions have been extensively studied from various perspectives, including computations for specific graded Lie algebras [3, 7, 11, 12, 25, 26, 31, 37] and theoretical aspects [36, 48, 49]. The concept of graded contraction is flexible with respect to changes in the underlying category of algebras and thus can be easily transferred to other structures, such as superalgebras [38] and Jordan algebras [30]. For a review of the state of the art in graded contractions, we refer to the recent paper by Draper, Sánchez-Ortega and Meyer [12].
@@
For generic graded contractions, which are studied in [36, 47, 48] and are also the subject of the present paper, the defining conditions are chosen in such a way that the contraction can be applied uniformly to all $G$-graded Lie algebras. This allows us to study such contractions from various perspectives: group cohomology, affine algebraic geometry and monoidal categories. As a result, we not only generalize and reinterpret existing results on classification of such contractions from [36, 48], but also place them into a wider context.
@@
The structure of the paper is as follows. In Section 2, we briefly review the necessary background on cohomology of abelian groups as well as gradings, contractions, degenerations, and graded contractions.
@@
In Section 3, we discuss the classification of generic graded contractions. To any subset $S$ of the set of unordered pairs in $G$, we associate the free abelian groups of “surviving defining relations” and “surviving higher-order identities” and use them to show that contractions with support $S$ are classified by a certain abelian group ${\rm H}^{2}_{S}(\mathbb{F}^{\times})$, which can be computed using the exact sequences presented in Theorem 19. In particular, for an algebraically closed or a real closed field $\mathbb{F}$, we reinterpret the results of Weimar–Woods [48].
@@
Section 4 is devoted to the affine variety of generic graded contractions. We study the action of an algebraic torus on this variety and the orbit closures of contractions. In Theorem 27, we describe the vanishing ideal of such an orbit closure. We show that the contractions in the closure of the unique open orbit are precisely the “continuous contractions” in the sense of [36] and relate them to Inönü–Wigner contractions using Theorem 19 and [47, Theorem VI.2].
@@
In Section 5, we discuss the Weimar–Woods conjecture that two graded contractions are isomorphic if and only if they are equivalent via normalization. We show that a generic $G$-graded contraction endows the identity endofunctor of the monoidal category of $G$-graded vector spaces with a lax monoidal structure. This allows us to prove a functorial version of the Weimar–Woods conjecture: two graded contractions define naturally isomorphic lax monoidal endofunctors if and only if they are equivalent via normalization (Proposition 36).
Finally, in Section 6, we summarize the results of the paper and discuss a few possible directions for future work.
@@ 2 Preliminaries
In this section, we review some basic facts about cohomology and extensions of abelian groups, see, e.g., [33, Chapter I] and [45, Chapter 6], contractions [39], gradings [17] and graded contractions of Lie algebras [12, 38].
2.1 Cohomology and extensions of abelian groups
Let $R$ be an associative unital ring and let $A$ and $B$ be $R$-modules. An extension of $A$ by $B$ is an $R$-module $E$ containing a submodule isomorphic to $B$ such that $E/B\simeq A$. In other words, the modules $A$, $B$ and $E$ fit into the short exact sequence
B{B}E{E}A,{A,}
[…]
$B{B}E{E^{\prime}}A{A}$$\scriptstyle{\varphi}$
is commutative. It follows that $\varphi$ is an isomorphism of $R$-modules, and this gives rise to an equivalence relation on the set of extensions. The set of equivalence classes of extensions of $A$ by $B$ is denoted by ${\rm Ext}_{R}(A,B)$ or, simply, ${\rm Ext}(A,B)$ when $R$ is clear from the context.
For two extensions $E$ and $E^{\prime}$ of $A$ by $B$ one can form the so-called Baer’s sum $E+E^{\prime}$ of extensions [33, Chapter III.2], which defines a binary operation on the set ${\rm Ext}(A,B)$ that makes it an abelian group [33, Chapter III.2, Theorem 2.1]. The neutral element of this group is the equivalence class of the direct sum $A\oplus B$; the elements of this class are called split extensions. In particular, all extensions are split when $A$ is projective or $B$ is injective as $R$-modules, [45, Sections 2.2 and 2.3]. In the case of abelian groups, i.e., $R=\mathbb{Z}$, this means that $A$ is free or $B$ is divisible.
Let $A$ and $B$ be abelian groups. The group ${\rm Ext}(A,B)$ is isomorphic to the symmetric part of the second cohomology group ${\rm H}^{2}(A,B)$ of $A$ with coefficients in $B$, considered as an $A$-module with the trivial $A$-action, see [45, Theorem 6.6.3]. More specifically, the abelian group ${\rm H}^{2}_{\rm sym}(A,B)$ is the quotient of the group of symmetric two-cocycles ${\rm Z}^{2}_{\rm sym}(A,B)$ of $A$ with values in $B$ by the group of coboundaries ${\rm B}^{2}(A,B)$,
${\rm H}^{2}_{\rm sym}(A,B):={\rm Z}^{2}_{\rm sym}(A,B)/{\rm B}^{2}(A,B).$
A symmetric two-cocycle $\varepsilon$ is a map from $A\times A$ to $B$ satisfying $\varepsilon(a_{1},a_{2})=\varepsilon(a_{2},a_{1})$ and
$\varepsilon(a_{1},a_{2})\varepsilon(a_{1}a_{2},a_{3})=\varepsilon(a_{1},a_{2}a_{3})\varepsilon(a_{2},a_{3})\quad\text{for all}\quad a_{1},a_{2},a_{3}\in A,$
where the group operations in $A$ and $B$ are written multiplicatively. For an arbitrary map $\alpha\colon A\to B$, its coboundary $d\alpha$ is the symmetric two-cocycle given by
$d\alpha(a_{1},a_{2}):=\alpha(a_{1})\alpha(a_{2})\alpha(a_{1}a_{2})^{-1}\quad\text{for all}\quad a_{1},a_{2}\in A.$
The correspondence between the groups ${\rm Ext}(A,B)$ and ${\rm H}^{2}_{\rm sym}(A,B)$ is constructed as follows. For a group $E$ in a short exact sequence $B\stackrel{{\scriptstyle\raisebox{-3.61664pt}[0.0pt][0.0pt]{\scriptsize$\iota$}}}{{\hookrightarrow}}E\stackrel{{\scriptstyle\raisebox{-3.61664pt}[0.0pt][0.0pt]{\scriptsize$\pi$}}}{{\twoheadrightarrow}}A$, consider a set-theoretic section $\sigma\colon A\to E$ of $\pi$. Then the element $\sigma(a_{1})\sigma(a_{2})\sigma(a_{1}a_{2})^{-1}$ belongs to $\ker\pi$, thus by exactness of the sequence there exists a unique $\varepsilon(a_{1},a_{2})\in B$ such that $\iota(\varepsilon(a_{1},a_{2}))=\sigma(a_{1})\sigma(a_{2})\sigma(a_{1}a_{2})^{-1}$ for all $a_{1},a_{2}\in A$. It is easy to verify that the map $\varepsilon\colon A\times A\to B$ is a two-cocycle, which is symmetric if and only if $E$ is abelian. The equivalence class $\varepsilon]\in{\rm H}^{2}(A,B)$ does not depend on the choice of the section $\sigma$.
A common approach to computing ${\rm Ext}(A,B)$ relies on finding a projective resolution of $A$ and then computing ${\rm Ext}(A,B)$ as the first (left) derived functor of the functor ${\rm Hom}(\cdot,B)$, see, e.g., [33, Chapter III]. In the case $R=\mathbb{Z}$, all subsequent derived functors are trivial. The choice of the projective resolution does not affect the group ${\rm Ext}(A,B)$ and can be adapted to the group $A$, which will be useful in Section 3.
@@ 2.2 Contractions and degenerations of Lie algebras
Contractions. Let $\mathfrak{L}:=(V,[\cdot,\cdot])$ be a real or complex Lie algebra with an underlying vector space $V$ and a Lie bracket $\cdot$$\cdot$]. Consider a continuous function $U\colon(0,1]\to{\rm GL}(V)$. A parameterized family of new Lie brackets on $V$ is defined in terms of the Lie bracket $\cdot$$\cdot$] in the following way:
$\forall t\in(0,1],\ \ \forall x,y\in V\colon\quad[x,y]_{t}:=U_{t}^{-1}[U_{t}x,U_{t}y].$ (1)
@@
If the limit $x,y]_{0}:=\lim_{t\to 0^{+}}[x,y]_{t}$ exists for any $x,y\in V$, then $\cdot,\cdot]_{0}$ is a well-defined Lie bracket on $V$ and the Lie algebra $\mathfrak{L}_{0}:=(V,[\cdot,\cdot]_{0})$ is called a one-parametric continuous contraction (or simply contraction) of the Lie algebra $\mathfrak{L}=(V,[\cdot,\cdot])$.
@@
Choosing a basis in each subspace $V^{(s)}$, we represent $U_{t}$ by a diagonal matrix with entries $t^{n_{s}}$, each repeated $\dim V^{(s)}$ times. Then the limit $\cdot,\cdot]_{0}$ exists if and only if the following constraints are satisfied:
@@
The constraints (3) are necessary and sufficient to ensure that $\mathfrak{L}(n),\mathfrak{L}(m)]\subset\mathfrak{L}(n+m)$. The associated graded algebra $\operatorname{gr}\mathfrak{L}$ of $\mathfrak{L}$ with respect to the above filtration is isomorphic to the contracted algebra $\mathfrak{L}_{0}$ under the contraction matrix $U(\varepsilon)$ from (2). This follows directly from the description of the structure tensor $\tilde{c}_{ij}^{k}$ of $\mathfrak{L}_{0}$, see, e.g., [39, Section 3]:
@@
When $\mathbb{F}$ is algebraically closed, the orbit $\mathcal{O}(\mu)$ is open in its Zariski closure $\overline{\mathcal{O}(\mu)}$ by the closed orbit lemma in [4]. Hence, the boundary $\overline{\mathcal{O}(\mu)}\setminus\mathcal{O}(\mu)$ is a union of orbits of strictly lower dimension. Moreover, all points of $\mathcal{O}(\mu)$ are regular in $\overline{\mathcal{O}(\mu)}$. Finally, the orbit $\mathcal{O}(\mu)$ is a constructible set, so in the case $\mathbb{F}=\mathbb{C}$, the closure of $\mathcal{O}(\mu)$ in Zariski topology coincides with its closure in the metric topology, see [35, Corollary 1, p. 60]. Therefore, over $\mathbb{C}$, the concepts of contraction and degeneration coincide.
@@ 2. $(ii^{\prime})$
To prove the necessity, it suffices to find a $G$-graded Lie algebra $\mathfrak{L}$ satisfying $(i)$ $\mathfrak{L}_{g},\mathfrak{L}_{h}]\neq 0$ for all distinct $g,h\in G$, $(ii)$ for all $g,h,k\in G$ there exist $x\in\mathfrak{L}_{g}$, $y\in\mathfrak{L}_{h}$, $z\in\mathfrak{L}_{k}$ such that $[x,[y,z]]$ and $[y,[z,x]]$ are linearly independent. For instance, consider the free Lie algebra $\mathfrak{F}$ on the generators $x_{g}$, $y_{g}$ and $z_{g}$, where $g$ runs through $G$. This algebra has a natural grading by setting the $G$-degree of generators $x_{g}$, $y_{g}$ and $z_{g}$ to be $g$. Since $x_{g},x_{h}]\neq 0$ for all $g\neq h$, condition $(i)$ implies $(i^{\prime})$ for the graded Lie algebra $\mathfrak{F}$. Since the commutators $x_{g},[y_{h},z_{k}]]$ and $y_{h},[z_{k},x_{g}]]$ are linearly independent in $\mathfrak{F}$, condition $(ii)$ implies $(ii^{\prime})$. ∎
@@
When $\varepsilon$ does not take value 0, conditions $(i^{\prime})$ and $(ii^{\prime})$ say that $\varepsilon$ is a symmetric two-cocycle with values in $\mathbb{F}^{\times}$ with trivial $G$-action: $\varepsilon\in{\rm Z}^{2}_{\rm sym}(G,\mathbb{F}^{\times})$, see Section 2.1.
@@
The obvious examples of the graded contractions are the constant maps $\varepsilon=1$ and $\varepsilon=0$. Their corresponding contracted algebras are respectively the original algebra $\mathfrak{L}$ and the abelian Lie algebra. Both these maps are examples of trivial graded contractions, that is, the graded contractions $\varepsilon\colon G\times G\to\mathbb{F}$ such that the for all $G$-graded Lie algebras $\mathfrak{L}$ the contracted algebra $\mathfrak{L}^{\varepsilon}$ is isomorphic to $\mathfrak{L}$ or to the abelian Lie algebra on the vector space $\mathfrak{L}$, see [48, Definition 2.6].
A more interesting example of a trivial graded contraction is that arising from the normalization of homogeneous components of the grading $\Gamma$. Let $\alpha\colon G\to\mathbb{F}^{\times}$ be an arbitrary map, then the map $d\alpha\colon G\times G\to\mathbb{F}^{\times}$ defined in Section 2.1 is a graded contraction of $\mathfrak{L}$. It is clear that $\mathfrak{L}^{d\alpha}$ is isomorphic to $\mathfrak{L}$. The choice of the notation $d\alpha$ for this kind of contractions is justified by the fact that it is the coboundary of the map $\alpha$ if we consider $\mathbb{F}^{\times}$ as a $G$-module with trivial action. Furthermore, when $\alpha$ is a group homomorphism then $d\alpha=1$.
Graded degenerations. Let $G$ be an abelian group. Consider a $G$-grading $\Gamma$ on a finite-dimensional vector space $V$ over an algebraically closed field $\mathbb{F}$, $V=\bigoplus_{g\in G}V_{g}$, with $\dim V_{g}=n_{g}\in\mathbb{N}\cup\{0\}$. Denote $\underline{n}=(n_{g})_{g\in G}$ and let $\mathcal{L}_{\underline{n}}(\mathbb{F})$ be the variety of Lie algebra laws on $V$ that make it a $G$-graded algebra. The set $\mathcal{L}_{\underline{n}}(\mathbb{F})$ is an algebraic set in the affine space $\wedge^{2}V^{*}\otimes V$, which is invariant under the natural action of the subgroup $\prod_{g\in G}{\rm GL}(V_{g})$ of ${\rm GL}(V)$. We will consider the subgroup $D(\Gamma)$ of $\prod_{g\in G}{\rm GL}(V_{g})$ consisting of the operators that act as (nonzero) scalars on each $V_{g}$. It is clear that $D(\Gamma)$ is an algebraic torus of dimension $|{\rm Supp}(\Gamma)|$.
@@
Consider the free Lie algebra $\mathfrak{F}$ on the generators $x_{g}$ and $y_{g}$, where $g$ runs through $G$. We can define a $G$-grading on $\mathfrak{F}$ by setting the $G$-degree of generators $x_{g}$ and $y_{g}$ to be $g$. Given a set $S\subset\mathcal{P}_{G}$, consider the ideal $I_{S}$ of $\mathfrak{F}$ generated by $\mathfrak{F}^{3}$ (the third term of the lower central series of $\mathfrak{F}$) and the commutators $\mathfrak{F}_{g},\mathfrak{F}_{h}]$ for all $\{g,h\}\notin S$. Then the quotient $\mathfrak{L}:=\mathfrak{F}/I_{S}$ is a $G$-graded Lie algebra, whose second support is $S$. Moreover, if $G$ is finite, the algebra $\mathfrak{L}$ is finite-dimensional. ∎
@@
Equivalences of graded contractions. Several natural equivalence relations can be introduced on the set of graded contractions. For the generic setting, some of them already appeared in [36] and [47]. There are also versions of these equivalences for a specific $G$-graded Lie algebra $\mathfrak{L}$.
Recall that, for $G$-graded algebras $A$ and $B$, an algebra isomorphism $\varphi\colon A\to B$ is an equivalence if it maps each homogeneous component of $A$ onto a homogeneous component of $B$: for any $g\in G$ there is $h\in G$ such that $\varphi\left(A_{g}\right)=B_{h}$, and $\varphi$ is a graded isomorphism if $\varphi\left(A_{g}\right)=B_{g}$ for all $g\in G$.
Definition 9 ([12]).
For a particular $G$-graded Lie algebra $\mathfrak{L}$, two graded contractions $\varepsilon$ and $\varepsilon^{\prime}$ are called equivalent (resp. strongly equivalent) if the contracted Lie algebras $\mathfrak{L}^{\varepsilon}$ and $\mathfrak{L}^{\varepsilon^{\prime}}$ are equivalent (resp. graded isomorphic).
The following is an analogue of strong equivalence for generic graded contractions: